OpenStudy (anonymous):

you have 276 feet of fencing to enclose a rectangular region. what is the maximum area that?

OpenStudy (anonymous):

here is a handy fact. The maximum occurs when it's a square

OpenStudy (anonymous):

Let x & y be the sides of the rectangle. perimeter 2x+2y=276 x+y=138 y=138-x area A=x*y=x(138-x)=138x-x^2 \[\frac{ dA }{ dx }=138-2x,\frac{ dA }{ dx }=0 ~gives138-2x=0,x=69\] \[at~x=69,\frac{ d^2A }{ dx^2 }= -2 ~<0,\] there is a maxima at x=69 Hence A is maximum when x =69 find y